OpenStudy (fairygirl0404):

What is the best way to solve the following equations: a. 5x^2+12x-3=0 b. x^2-4x=8 c. 4x^2-25=0 d. x^2-5x+6=0 solve by: quadratic formula, completing the square, square root method, factoring

OpenStudy (fairygirl0404):

I think they are... a. quadratic formula b. square root method c. factoring d. completing the square Are these correct?

Directrix (directrix):

d. x^2-5x+6=0 factors (x -3)*(x - 2)

Directrix (directrix):

c. 4x^2-25=0 will factor but taking square roots may be easier. 4x^2 = 25 2x = +5 or 2x = -5 and so on

Directrix (directrix):

x^2-4x=8 Maybe complete the square 5x^2+12x-3=0 Maybe the Quadratic Formula I'm not saying that any of my thoughts are correct. I would factor those which will factor and QF the rest. But, that is not the question.

Directrix (directrix):

@fairygirl0404 What happened to our other helpers. We need more input.

Directrix (directrix):

@agent0smith

OpenStudy (fairygirl0404):

@Directrix I agree with you about factoring and QF the rest...you are right it is asking to break them up more. Thanks for you input. Greatly apperciated. :)

Directrix (directrix):

You are welcome.

OpenStudy (fairygirl0404):

@Directrix because I voted yur response Best, does that mean no one will see my problem now? Or is still open till I close it?

Directrix (directrix):

Here's my list: 1. Quadratic Formula 2. Completing the Square 3.Square Roots 4. Factoring.

OpenStudy (agent0smith):

a. 5x^2+12x-3=0 - QF b. x^2-4x=8 - QF or complete the square c. 4x^2-25=0 - factor d. x^2-5x+6=0 - factor

Directrix (directrix):

If you close the problem helpers can still work on it. And, there can be more than one Best Response." Closing means the question goes over into another folder on OS but we can still see it.

OpenStudy (fairygirl0404):

@Directrix LOL I'll let you know the outcome when I get my paper back. Thanks again for all your help.

Directrix (directrix):

Check out @agent0smith 's post.

Directrix (directrix):

Alrighty, then, @fairygirl0404 close if you are ready.

OpenStudy (fairygirl0404):

I did but there can't be two factors. It is a = 1 and so on.

OpenStudy (agent0smith):

...What?